In a large? city, 65 ?% of people pass the? drivers' road test. Suppose that every? day, 300 people independently take the test. In a large? city,65?% of people pass the? drivers' road test. Suppose that every? day, 300 people independently take the test.

After a great many? days, according to the Empirical? Rule, on about? 95% of these? days, the number of people passing will be as low as? _____ and as high as? _____. (Hint: Find two standard deviations below and two standard deviations above the? mean.)
Answer #1

Let X denotes the number of people passing out of 300 people who independently take the test everyday.

X ~ Binomial(300, 0.65)

The probability mass function of X is

P(X=x) = \binom{300}{x}*0.65^x*(1-0.65)^{300-x},x=0,1,2,...,300

Here

Mean E(X) = 300*0.65 = 195

Standard deviation sd(X) = \sqrt{300*0.65*(1-0.65)} = 8.26136

Mean - (2*sd)  = 195 - (2* 8.26136) = 178.477

Mean + (2*sd)  = 195 + (2* 8.26136) = 211.523